OpenStudy (anonymous):

find the x-intercepts of the parabola with vertex (-3,-14) and y intercept (0,13). Write answer in (x1,y1),(x2,y2) form. Round to the nearest hundredth.

OpenStudy (watchmath):

From the vertex information, we know that the parabola is of the form\ \(y=a(x+3)^2-14\) Plug in (0,13) we have \(13=9a-14\) \(9a=27\) \(a=3\) So the parabola is \(y=3(x+3)^2-14=3x^2+18x+13\) Setting \(3x^2+18x+13=0\) Solve the x by using the quadratic formula.